OpenStudy (rayford):

2x+y=6 -4x+2y=4 solve using the elimination method

OpenStudy (anonymous):

2(2x+y)=2(6) 4x+2y=12 add them -4x+2y=4 -4x+4x+2y+2y=16 4y=16

OpenStudy (anonymous):

elimination mean's solve for one variable. this can be done by simple subtraction of on equation from the other.

OpenStudy (anonymous):

divide both sides by 4 and you got your y. Use that y and plug it in into any equation and then solve for x

OpenStudy (rayford):

thank youits! 1,4!

OpenStudy (rayford):

may you answer another?

OpenStudy (anonymous):

good job mate!

OpenStudy (anonymous):

Yeah just post it in another question

OpenStudy (rayford):

x+2y=10 x-y=4 use substituion